OpenStudy (anonymous):

a box of chocolates contains 12 chocolates with nuts and 24 24 with out nuts. Determine the probability of choosing 2 chocolates with nuts in a row without replacement

OpenStudy (anonymous):

P( first choco with nut) = 12/ 36 = 1/3 P ( second with nut) = 11/35 => P ( both) = 11/ 45